OpenStudy (anonymous):

How do I solve this? Simplify. (x+1) (x-5)= Is the answer -4x?

OpenStudy (anonymous):

x2-5x+x-5 x2-4x-5

OpenStudy (anonymous):

i got x^2-4x-5

OpenStudy (anonymous):

No, (x+1) (x-5)= x^2+6x-5

OpenStudy (anonymous):

well can you walk me through the steps?

OpenStudy (anonymous):

check your work @Sjano

OpenStudy (anonymous):

trick to remebering. first outside inner last

OpenStudy (anonymous):

well using the foil can you show me what steps are for each letter?

OpenStudy (anonymous):

(x+1) (x-5)=(x*x)+(x*-5)+(1*x)+(1*-5)=x^2-4x-5 @kiamonstaa thanks

OpenStudy (anonymous):

first multiple the x's to get x2 thenmultipe x and -5 toget -5x then x times 1 to get 1 x then -5 times 1 to get -5 x2-5x+x-5 x2-4x-5

OpenStudy (anonymous):

welcome!

OpenStudy (anonymous):

If you use foil, wouldn't the L standing for last be 1-5? being -4?

OpenStudy (anonymous):

multiply 1 and -5

OpenStudy (anonymous):

you still get the same answer